emilybethch:

A tub filled with 50 quarts of water empties at a rate of 2.5 quarts per minute. Let w = quarts of water left in the tub and t = time in minutes. A 2-column table with 3 rows. The first column is labeled time (t) with entries 0, 2, 4. The second column is labeled quarts of water (W) with entries 50, 45, 40. Choose the correct answers. Which equation models the relationship? Is there a viable solution when time is 30 minutes? HELP PLS!!!!

Hoodmemes:

Where are the answers ?

emilybethch:

t=2.5-50w t=50-2.5w w=2.5-50t w=50-2.5t

emilybethch:

and....

emilybethch:

no, time will stop at 30 min. no, he tub will be empty then. yes, the tub will contain 25 quarts of water. yes, time continues forever.

jhonyy9:

so bc the tub filled with 50 quarts of water and given that 2,5 quarts / min. from this result what ? how many get in 10 minutes ?

jhonyy9:

emily please collaborate

jhonyy9:

2,5 quarts in 1 minute x quarts in 10 min. x = ?

jhonyy9:

cross multiplie and calcule the x

jhonyy9:

x*1 = 2,5 *10 x = ?

jhonyy9:

x = 25 quarts so from this result that the tub filled till the half part in 10 minutes yes ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 in how much minutes is the tub filled with 50 quarts ? \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

25 quarts in 10 min. 50 quarts in x min. x = (50*10)/25 = ?

emilybethch:

um let me see

emilybethch:

1 attachment
jhonyy9:



emilybethch:

A tub filled with 50 quarts of water empties at a rate of 2.5 quarts per minute. Let w = quarts of water left in the tub and t = time in minutes.

emilybethch:

yeh

jhonyy9:

Which equation models the relationship?

jhonyy9:

so from these above wrote this is very easy to get it

jhonyy9:

25 quarts in 10 min. 50 quarts in x min. x = (50*10)/25 = ? what result from this than you ve get the value of x ?

emilybethch:

i still dont understand math or whatever this devilish thing is

jhonyy9:

emily please little attention !!! we ve got above that from this tub in 10 minutes are empty 25 quarts of water and you know that the tub contain 50 quarts of water how you think in the tube how much water will be after 20 minutes ?

jhonyy9:

do you understand please this above wrote ?

jhonyy9:

in 10 min. empty 25 quarts of water in 2*10 min. empty 2*25 quarts of water yes ?

jhonyy9:

so from this result that after 20 minutes the tube will be empty full ,with no water yes ?

emilybethch:

i think so

jhonyy9:

and now you can answer this easy Is there a viable solution when time is 30 minutes?

emilybethch:

kinda whats a viable?

jhonyy9:

one possible

emilybethch:

idk

emilybethch:

i is confused

jhonyy9:

why ? we ve get above that the tube have nothing water after 20 minutes - yes ? so what about this after 30 minutes ?

emilybethch:

oh um....

jhonyy9:

ok ? do you understand it now sure clearly please ?

emilybethch:

it has nothing?????

jhonyy9:

exactly

emilybethch:

i think i do

jhonyy9:

so there is a viable solution after 30 minutes ?

emilybethch:

yes?????

emilybethch:

i feel like im wrong

jhonyy9:

NNNNOOOOOOO bc. after 20 minutes the tub has nothing inside

emilybethch:

ooooooooooooh!!!!!

emilybethch:

i get it now

emilybethch:

i thnk.....

jhonyy9:

now is clearly sure

emilybethch:

uh huh

jhonyy9:

do you have again questions on this ?

emilybethch:

so what about the other question????

emilybethch:

about t=2.5-50w t=50-2.5w w=2.5-50t w=50-2.5t

jhonyy9:

so on what you think ?

jhonyy9:

ok

jhonyy9:

a little logic from there how you think that is correct ?

jhonyy9:

we ve solved it above

emilybethch:

what u mean????

jhonyy9:

please check them step by step bc. the water is empty so the equation need begin with water yes ? so w

emilybethch:

yeh

jhonyy9:

inside the tub there is 50 quarts of water so from this result that how much will be empty from this so this quantity need being subtracted from 50 yes ?

emilybethch:

yeh

jhonyy9:

ok so now please check these equations and try get where you see that begin with w = 50 -

emilybethch:

okay

jhonyy9:

you dont see it there ?

emilybethch:

uh huh

emilybethch:

i get it now thank you, budy!!

jhonyy9:

really sure ?

emilybethch:

i think so....

jhonyy9:

what is the correct equation ?

emilybethch:

its w=50-2.5t no, the tub will be empty by then ????

jhonyy9:

GREAT hope helped you understandably

emilybethch:

thank you so much!!

jhonyy9:

was my pleasure and any time

emilybethch:

:)